OpenStudy (anonymous):

Solve the equation for x. Round your answer to two decimal places. Do not include "x =" in your answer. 6 ex = 25

OpenStudy (anonymous):

introduce \[\ln \]to both sides of your equation

OpenStudy (anonymous):

So that you have \[6\ln ex=\ln25\] and remember \[blna= lna^b\]

OpenStudy (anonymous):

introduce \[\ln \]to both sides of your equation

OpenStudy (anonymous):

okayy;}

OpenStudy (anonymous):

1.43?

OpenStudy (anonymous):

Try again

OpenStudy (anonymous):

-.-no it was correct.. thanks for the help tho